# Automated Defect Detection for Mass-Produced Electronic Components Based on YOLO Object Detection Models

### v1

@@ Abstract
Deep learning (DL), You Only Look Once (YOLO), automated defect detection, camera sensor, automated optical inspection (AOI), dual inline package (DIP) switch, ConSinGAN.
@@ I Introduction
Over the past decade, DL [3, 4, 5, 6] has been widely applied to various fields, including image recognition, natural language processing, autonomous driving, and object classification problems. Further, DL has also been used for medical radiography image detection and classification [7, 8, 9, 10], A. Jo and W. Lee, [11] proposed a method of deep learning models for material discrimination and quantitation. S. M. J. Jalali et al., [12] proposed a method of deep reinforcement learning-based ensemble algorithm that integrates optimized deep learning models to minimize wind power forecasting errors for two wind power datasets. In [13], a method of deep learning was proposed by E. Jeong et al. for heterogeneous inference parallelization.
@@ III-A2 ConSinGAN
III-B YOLO Learning Series
YOLO is adopted as a DL-based object detection algorithm. Compared with traditional object detection algorithms, the YOLO model can quickly perform object detection and classification within a single neural network. The YOLO architecture is a convolutional neural network (CNN) that includes multiple convolutional and fully connected layers, segments an image into S × S grids, and then detects multiple objects in each grid. In addition, the selection criteria of the different YOLO versions for DIP testing were based on their floating-point operations (FLOPs) and mAP. To compare different YOLO versions, we list the performance in terms of FLOPs, and of YOLO v3, v4, v5, v6, v7, v8, and v9, as shown in Table IV. YOLOv7 [29] is chosen as the proposed DIP testing model. In our simulation, YOLOv3 [27], YOLOv4  [28], and the current newest YOLOv9 [33] are considered for comparison. As shown in Table VI, the YOLOv7 slightly outperforms YOLOv9 in terms of accuracy but is slightly slower. YOLOv5 [30] and YOLOv6 [31] have similar FLOP performance to YOLOv7, but their mAP is less than that of YOLOv7. YOLOv8 has higher computation complexity than YOLOv7 in terms of FLOP by 63%; however, the mAP improved by only 0.1%. Therefore, YOLOv5, v6, and v8 are excluded from our main simulations.
@@ III-B2 YOLOv4 [28]
First, YOLOv4 uses CSPDarknet53 to establish the primary backbone neural network, making it a lightweight model while maintaining high accuracy. Second, YOLOv4 adopts spatial pyramid pooling [54] and path aggregation network [55] in the neck section to combine feature maps at different scales and detect different sizes of objects. Finally, the YOLOv4 employs CIoU-loss to evaluate the distance between the targeted and the predicted boxes and improve the accuracy.
@@ III-B3 YOLOv7 [29]
In Fig. 9, YOLOv7 uses extended ELAN (E-ELAN), which develops the original ELAN architecture through expanding, shuffling, and merging. These approaches allow the network to learn more diverse features without disrupting the original gradient paths. The head part of YOLOv7, which is used for object detection and localization, continues to utilize the head network structure of YOLOR [56]. The CBS block includes the convolutional layer, batch normalization, and the Silu activation function. The CBM block consists of the convolutional layer, batch normalization, and the Sigmoid activation function. REP block represents the RepVGG. The difference between MP-1 and MP-2 is the output size within the block. The difference between E-ELAN and ELAN-W is the output size within the block. The main improvements of YOLOv7 are as follows:
  • In RepConv, the identity connection (id) disrupts the residual in ResNet and the concatenation in DenseNet. However, RepConv needs to retain features that provide greater gradient diversity for different feature maps. A novel weight parameterization approach is proposed in [57] for RepConv to replace id. In Fig. 10, the planned reparametrized convolution designed for ResNet is called RepConvN [29]. Since PlainNet in Fig. 10(a) inherently lacks residual or concatenation characteristics, RepConv in Fig. 10(b) can be applied directly. Since ResNet in Fig. 10(c) inherently has residual characteristics, RepConv in Fig. 10(d) leads to a decrease in accuracy. Based on the above reasons, YOLOv7 uses RepConvN (RepConv without id) to design a network architecture.
@@ III-B4 YOLOv9 [33]
  • GELAN combines the functions of CSPNet and ELAN networks. It splits the input of each computational unit into two parts: one part is directly passed to the output, while the other part is transformed by the computational unit. Then, these two parts of computation are combined to form the final output. Thus, GELAN integrates information between computational units while enhancing flexibility and generalization.
@@ IV-B Performances of YOLO with/without ConSinGAN
The YOLO model is trained for each aspect of the generated image data. These 3,183 images are divided into two categories: surface defects (surface overflow, surface scratches, and surface dirt) and pin defects (pin misalignment). We compare YOLOv3, v4, v7, and v9 with/without the ConSinGAN augmented sets to identify the optimal model. The training set is conducted on 80% of the dataset, i.e., 2,547 images. The validation set is adopted on the other 10% (318 images). The remaining 10% (318 images) are used for the testing set. Within this dataset, we define two distinct defect types that include: 1) Surface defects account for approximately 50.7% of defects, and 2) Misaligned pins account for 49.3%. Table VI demonstrates the performances of YOLOv3, v4, v7, and v9 with/without the ConSinGAN in terms of average precision (mAP) rate, $PRE$, RECREC, $F1$, FPRFPR, and $TNR$ evaluations of all six YOLO models.

### v2

@@ Abstract
Automated defect detection, automated optical inspection (AOI), camera sensor, ConSinGAN, deep learning (DL), dual in-line package (DIP) switch, you only look once (YOLO).
@@ I Introduction
Over the past decade, DL [3, 4, 5, 6] has been widely applied to various fields, including image recognition, natural language processing, autonomous driving, and object classification problems. Further, DL has also been used for medical radiography image detection and classification [7, 8, 9, 10], A. Jo and W. Lee, [11] proposed a DL-based model for material discrimination and quantitation. S. M. J. Jalali et al., [12] proposed the deep reinforcement learning-based ensemble algorithm that integrates optimized deep learning models to minimize wind power forecasting errors for two wind power datasets. In [13], a DL-based model was proposed by E. Jeong et al. for heterogeneous inference parallelization.
@@ III-A2 ConSinGAN
III-B YOLO Object Detection Model Series
YOLO is adopted as a DL-based object detection algorithm. Compared with traditional object detection algorithms, the YOLO model can quickly perform object detection and classification within a single neural network. The YOLO architecture is a convolutional neural network (CNN) that includes multiple convolutional and fully connected layers, segments an image into S × S grids, and then detects multiple objects in each grid. In addition, the selection criteria of the different YOLO versions for DIP testing were based on their floating-point operations (FLOPs) and mAP. To compare different YOLO versions, we list the performance in terms of FLOPs, and of YOLO v3, v4, v5, v6, v7, v8, and v9, as shown in Table IV. YOLOv7 [29] is chosen as the proposed DIP testing model. In our simulation, YOLOv3 [27], YOLOv4  [28], and the current newest YOLOv9 [33] are considered for comparison. As shown in Table VI, the YOLOv7 slightly outperforms YOLOv9 in terms of accuracy but is slightly slower. YOLOv5 [30] and YOLOv6 [31] have similar performance to YOLOv7 in terms of FLOPs, but their mAP is less than that of YOLOv7. YOLOv8 has higher computation complexity than YOLOv7 in terms of FLOP by 63%; however, the mAP improved by only 0.1%. Therefore, YOLOv5, v6, and v8 are excluded from our main simulations.
@@ III-B2 YOLOv4 [28]
First, YOLOv4 uses CSPDarknet53 to establish the primary backbone neural network, making it a lightweight model while maintaining high accuracy. Second, YOLOv4 adopts spatial pyramid pooling (SPP) [54] and path aggregation network [55] in the neck section to combine feature maps at different scales and detect different sizes of objects. Finally, the YOLOv4 employs complete intersection over union (CIoU)-loss to evaluate the distance between the targeted and the predicted boxes and improve the accuracy.
@@ III-B3 YOLOv7 [29]
In Fig. 9, YOLOv7 uses an extended efficient layer aggregation network (E-ELAN), which develops the original ELAN architecture through expanding, shuffling, and merging. These approaches allow the network to learn more diverse features without disrupting the original gradient paths. The head part of YOLOv7, which is used for object detection and localization, continues to utilize the head network structure of YOLOR [56]. The convolution batch normalization silu (CBS) block includes the convolutional layer, batch normalization (BN), and the Mish activation function [57]. The spatial pyramid pooling cross-stage partial (SPPCSPC) is the first module in the head region, and it receives input from E-ELAN in the backbone. The SPPCSPC is composed of a cross-stage partial network (CSPNet) with an SPP block instead of a dense block as in YOLOv5 [30]. The CBM block consists of the convolutional layer, batch normalization, and the sigmoid-weighted linear unit (silu) activation function. The reparameterization (REP) block represents the RepVGG. The difference between max pooling (MP)-1 and MP-2 is the output size within the block. The difference between E-ELAN and ELAN-W is the output size within the block. The main improvements of YOLOv7 are as follows:
  • In re-parameterized refocusing convolution (RepConv), the identity connection (id) disrupts the residual in the residual neural network (ResNet) and the concatenation in dense convolutional network (DenseNet). However, RepConv needs to retain features that provide greater gradient diversity for different feature maps. A novel weight parameterization approach is proposed in [58] for RepConv to replace id. In Fig. 10, the planned reparametrized convolution designed for ResNet is called RepConvN (RepConv without identity connection (id)) [29]. Since PlainNet in Fig. 10(a) inherently lacks residual or concatenation characteristics, RepConv in Fig. 10(b) can be applied directly. Since ResNet in Fig. 10(c) inherently has residual characteristics, RepConv in Fig. 10(d) leads to a decrease in accuracy. Based on the above reasons, YOLOv7 uses RepConvN to design a network architecture.
@@ III-B4 YOLOv9 [33]
  • GELAN combines the functions of the CSPNet and ELAN networks. It splits the input of each computational unit into two parts: one part is directly passed to the output, while the other part is transformed by the computational unit. Then, these two parts of computation are combined to form the final output. Thus, GELAN integrates information between computational units while enhancing flexibility and generalization.
@@ IV-B Performances of YOLO with/without ConSinGAN
The YOLO model is trained for each aspect of the generated image data. These 3,183 images are divided into two categories: surface defects (surface overflow, surface scratches, and surface dirt) and pin defects (pin misalignment). We compare YOLOv3, v4, v7, and v9 with/without the ConSinGAN augmented sets to identify the optimal model. The training set is conducted on 80% of the dataset, i.e., 2,547 images. The validation set is adopted on the other 10% (318 images). The remaining 10% (318 images) are used for the testing set. Within this dataset, we define two distinct defect types that include: 1) Surface defects account for approximately 50.7% of defects, and 2) Misaligned pins account for 49.3%. Table VI demonstrates the performances of YOLOv3, v4, v7, and v9 with/without the ConSinGAN in terms of mAP rate, $PRE$, RECREC, $F1$, FPRFPR, and $TNR$ evaluations of all six YOLO models.